### v1

@@ Abstract
Collaborative mechanisms allow benchmarks to be updated continuously and adjust to the changing requirements and new use cases. This paradigm is employed for example in the field of machine learning, but up until now there were no examples of truly open and collaborative benchmarks for RDF systems. In this demo paper we present the collaboration functionalities of RiverBench, an open, multi-task RDF benchmark suite. Owing to its fully open and community-driven design, RiverBench allows any researcher or practitioner to submit a new dataset or benchmark task, report performed benchmark runs, and edit any resource in the suite. RiverBench’s collaboration system is itself largely based on RDF and Linked Data mechanisms, and every resource in the suite has machine-readable RDF metadata. The showcased functionalities together make up a first-of-a-kind fully open and collaborative RDF benchmark suite. These features are meant to encourage other researchers to contribute to RiverBench, and make it a long-term project sustained by the community.
@@ 1 Introduction
In this demo paper we showcase the recently-developed collaborative features of RiverBench [6], an open, multi-task benchmark suite for RDF systems. RiverBench goes beyond the presented state of the art in RDF benchmarks, by being designed to be fully open, transparent, and community-driven from the start. It allows submitting new benchmark tasks and datasets, reporting benchmark results, and editing any other aspect of the suite. The collaboration architecture is based on RDF metadata and Linked Data mechanisms, making every resource in RiverBench machine-accessible.
@@ 2 Collaborative Architecture of RiverBench
Benchmark datasets are added to the suite through a public community process, visualized in Figure 1. Any practitioner or researcher is invited to submit their dataset, as long as it meets basic technical requirements and can be published under a permissive license. Firstly, the contributor fills out the dataset proposal form on GitHub, giving details about the dataset’s source, license, original use case, etc. The proposal is then reviewed by an administrator, who checks if the dataset meets the publicly-available criteria for new datasets (e.g., open license, properly specified authorship, sufficient size, clear use case). The administrator then creates a new GitHub repository for the dataset, and instructs the contributor on how to proceed. The contributor then uploads the dataset for processing and fills out the metadata file in the Turtle language, describing the dataset in a structured manner. The dataset and the metadata are then automatically processed by CI/CD: validating its contents, re-packaging the dataset in several formats, adding more metadata, generating documentation pages, and publishing it on the website under a permanent URL (PURL).
@@
Proposing new benchmark tasks follows a very similar public community process. The contributor also fills out a proposal form on GitHub, which is reviewed by an administrator who checks if the task meets the pre-defined criteria. Afterwards, the contributor is asked to write down the description and metadata of the task in a Turtle file. The new task is then automatically processed by CI/CD (validation, packaging, documentation generation, publishing).
The source code of tasks, datasets, profiles (groups of datasets), documentation pages, and other resources of RiverBench are hosted on GitHub as Markdown or Turtle files in one of the several repositories. To make editing these resources as easy as possible, every page on the RiverBench website has an “Edit this page” button (visible at the top of Figure 2). The button immediately redirects the user to the correct file on GitHub, where they can edit it and submit a pull request with the changes. After the pull request is validated by a CI job and accepted by the maintainer, the changes automatically appear on the website via a CI/CD pipeline.
@@ 4 Conclusion
To the best of our knowledge, RiverBench is the first RDF benchmark suite to implement collaborative features so widely. We hope that these easy-to-use functionalities will encourage the Semantic Web and Knowledge Graph communities to contribute to RiverBench, ensuring its long-term relevance and usefulness. In contrast to the now-defunct Liquid Benchmarks platform [2], RiverBench is hosted on permanently free infrastructure, which ensures its longevity. We plan to continue expanding the suite’s documentation and automation system, to match the changing expectations of the community.

### v2

@@ Abstract
Collaborative mechanisms allow benchmarks to be updated continuously and adjust to the changing requirements and new use cases. This paradigm is employed for example in the field of machine learning, but up until now there were no examples of truly open and collaborative benchmarks for RDF systems. In this demo paper we present the collaboration functionalities of RiverBench, an open, multi-task RDF benchmark suite. Owing to its fully open and community-driven design, RiverBench allows any researcher or practitioner to submit a new dataset or benchmark task, report performed benchmark runs, and edit any resource in the suite. RiverBench’s collaboration system is itself based on RDF and Linked Data mechanisms, and every resource in the suite has machine-readable RDF metadata. The showcased functionalities together make up a first-of-a-kind fully open and collaborative RDF benchmark suite. These features are meant to encourage other researchers to contribute to RiverBench, and make it a long-term project sustained by the community.
@@ 1 Introduction
Contribution.
In this demo paper we showcase the recently-developed collaborative features of RiverBench [6], an open, multi-task benchmark suite for RDF systems. RiverBench is designed to be fully open, transparent, and community-driven from the start. It allows submitting new benchmark tasks and datasets, reporting benchmark results, and editing any other aspect of the suite. The collaboration architecture is based on RDF metadata and Linked Data mechanisms, making every resource in RiverBench machine-accessible. The presented approach goes beyond the current state of the art by being (to the best of our knowledge) the first RDF benchmark suite with such extensive open collaboration functionalities. It is also a major technical advancement over the aforementioned benchmark projects in other fields of computer science, which do not fully employ the FAIR principles, machine-readable RDF metadata, or the Linked Data mechanisms.
@@ 2 Collaborative Architecture of RiverBench
Benchmark datasets are added to the suite through a public community process, visualized in Figure 1. Any practitioner or researcher is invited to submit their dataset, as long as it meets basic technical requirements and can be published under a permissive license. Firstly, the contributor fills out the dataset proposal form on GitHub, giving details about the dataset’s source, license, original use case, etc. The proposal is then reviewed by RiverBench curators, who check if the dataset meets the publicly-available criteria for new datasets (e.g., open license, properly specified authorship, sufficient size, clear use case) and if it would be a valuable addition. The curators are members of the RDF benchmarking community, who besides reviewing dataset proposals also have a deciding voice in major decisions shaping RiverBench. As of November 2024, there are four registered curators from three different institutions [^1] https://w3id.org/riverbench/v/dev/documentation/maintainers. After the proposal is accepted by the curators, a technical administrator creates a new GitHub repository for the dataset, and instructs the contributor on how to proceed. The contributor then uploads the dataset for processing and fills out a metadata file in the Turtle language, describing the dataset in a structured manner. The dataset and the metadata are then automatically processed by CI/CD: validating its contents, re-packaging the dataset in several formats, adding more metadata, generating documentation pages, and publishing it on the website under a permanent URL (PURL).
@@
Proposing new benchmark tasks follows a very similar public community process. The contributor also fills out a proposal form on GitHub, which is reviewed by the curators who check if the task meets the pre-defined criteria. Afterwards, the contributor is asked to write down the description and metadata of the task in a Turtle file. The new task is then automatically processed by CI/CD (validation, packaging, documentation generation, publishing).
The source code of tasks, datasets, profiles (groups of datasets), documentation pages, and other resources of RiverBench are hosted on GitHub as Markdown or Turtle files in one of the several repositories. To make editing these resources as easy as possible, every page on the RiverBench website has an “Edit this page” button (visible at the top of Figure 2). The button immediately redirects the user to the correct file on GitHub, where they can edit it and submit a pull request with the changes. After the pull request is validated by a CI job and accepted by a curator or technical administrator, the changes automatically appear on the website via a CI/CD pipeline.
@@ 4 Conclusion
To the best of our knowledge, RiverBench is the first RDF benchmark suite to implement collaborative features so widely. We hope that these easy-to-use functionalities will encourage the Semantic Web and Knowledge Graph communities to contribute to RiverBench and join the project as curators, ensuring its long-term relevance and usefulness. The interest of the community ultimately depends on whether the project fulfills an important need or not – in this case, if it useful for benchmarking RDF systems. Currently, RiverBench can be used with nine diverse, scientifically relevant benchmark tasks, but this is expected to be expanded in the future. Based on the early feedback from other researchers, we plan to expand the suite’s infrastructure to support more use cases and tasks. We will also continue improving the suite’s documentation and automation system, to match the changing expectations of the community.
In contrast to the now-defunct Liquid Benchmarks platform [2], RiverBench is hosted on permanently free infrastructure (GitHub, w3id.org, and the nanopublication network), which should ensure its longevity. While w3id.org and the nanopublication network are maintained by multiple independent actors and therefore should be stable, GitHub could conceivably cease its operations in the future, or stop providing free services. RiverBench is prepared for that eventuality, as all its resources are hosted under permanent URLs, which can be easily redirected to a different service provider. Because the core features of the suite only require a static file hosting server, the minimal migration would not be problematic. The only code is executed as part of CI/CD pipelines, which could be ported to alternative services (e.g., GitLab).